031 10.0 points You want to enlarge a skating surface so you stand on the ice at one end and aim a hose horizontally to spray water on the schoolyard pavement. Water leaves the hose at 1.7 kg/s with a speed 41 m/s. If your mass is 62 kg, what is your recoil acceleration? Neglect friction and the mass of the hose. Answer in units of m/s2
Answer #1

Force due to water F = 1.7 × 41 kg.m/s^2

=) F = 69.7 N

Now acceleration of m=62kg is

a = F/m = 69.7/62 = 1.124 m/s^2 = 1.12m/s^2